OpenStudy (anonymous):

someone please solve -2-10/3 my answer is not the same as the books?

OpenStudy (anonymous):

i think the answer is -12/3 = -4

OpenStudy (sandra):

if this is (-2) - 10/3 , then the first thing you want to do is get everything with the same denominator

OpenStudy (sandra):

so -2 *3/3 = 6/3

OpenStudy (sandra):

so the problem is really -6/3 - 10/3

OpenStudy (sandra):

or -16/3

OpenStudy (sandra):

does that make sense?

OpenStudy (anonymous):

yes thats what i got then you have to divide by 4. so then multiply by the reciprocal and i got -16/12 which simplifies to -4/3 right?

OpenStudy (sandra):

hmmm, what do you mean divide by 4?

OpenStudy (anonymous):

well that was part of the equation. its x + 4z=-2 and x= 10/3

OpenStudy (sandra):

oh I did not see that part. so you're solving for z

OpenStudy (anonymous):

yeah i didnt include it in my question but yeah solving for z

OpenStudy (sandra):

so right 4z = -16/3

OpenStudy (sandra):

so to isolate z, you divide both sides by 4

OpenStudy (sandra):

so z = -16/(3*4) = -16/12 = -4/3

OpenStudy (sandra):

also, go ahead and plug it back in to the equation

OpenStudy (anonymous):

thats exactly what i had but the book said 1/3

OpenStudy (sandra):

since now you have both the x value, and the z value

OpenStudy (sandra):

put them back in and make sure the equation holds true

OpenStudy (sandra):

(even if you just use your calculator)

OpenStudy (anonymous):

the book said -1/3

OpenStudy (anonymous):

yeah ill try it

OpenStudy (sandra):

if you put them back in, then 10/3 + 4*(-4/3) = - 2

OpenStudy (sandra):

so 10/3 - 16/3 = -2

OpenStudy (sandra):

-6/3 = -2

OpenStudy (sandra):

-2 = -2

OpenStudy (sandra):

so you were right, back of the book was wrong :p

OpenStudy (anonymous):

lol okay good cause that was really irritating me i was sitting there thinking "how am i wrong?!" ahah

OpenStudy (sandra):

haha yeah =) it's good you questioned it :)

OpenStudy (anonymous):

thank you